# Belief Bias Identification

### v2

@@ 4.2 Complete model regressions
Variables in equations (14) and (15) are identical to the baseline model with the only exception of variable $c$, the relative measure of confirmation, as described in section 2.2.2. Equation (16) is almost identical to equation (11) in the interpretation of its variables and the $\nu$ parameter indicating over or underconfidence ($\nu>1$ vs $\nu<1$). The inclusion of an intercept $\eta$ and an error term $\varepsilon_{v}$ are the only differences. [^17] Similarly to the baseline model, equations (14) and (15) do not have an intercept for the very same reason as outlined in section 4.1.
Parameters $\rho$ and $\delta$ in equations (14) and (15) indicate confirmation/disconfirmation bias, and base rate neglect/overuse respectively. As in the baseline case, equations (14) and (15) differ by acknowledging that these parameters may differ between successes and failures ($\rho_{s}$ vs $\rho_{f}$) and ($\delta_{s}$ vs $\delta_{f}$). However, the notable difference between equations (14), (15) and (9),(10) lies in accommodating the biases of section 2.2.3. This involves the inclusion of three dummy variables – $IPref$, $ISeq_{s}$, and $ISeq_{f}$ – interacting with the number of successes (in (14)) or failures (in (15)). $IPref$ equals 1 when the subject faces a decision with preferences over the state-space (that is, when beliefs about a dollar urn where reported) and 0 otherwise. $ISeq_{s}$ ($ISeq_{f}$) equals 1 when the last three balls observed in the second sequence are red (blue). By doing so, one can distinguish motivated beliefs (optimism and pessimism), sequence-related biases (gambler’s fallacy and hot hand fallacy) and unrelated over/under inference. The presence of those biases is tested by comparing whether the estimated parameters are significantly different from their Bayesian values. Table 1 summarizes the baseline and complete models, and specifically shows the ranges of values of the different parameters, which correspond to each bias.
@@
$\gamma_{s}$ or $\gamma_{f}<1$: Underinference $\alpha_{0}$ or $\beta_{0}<1$: Underinference
Base Rate Biases
$\delta_{s}$ or $\delta_{f}>1$: Base Rate Overuse $\delta_{s}$ or $\delta_{f}>1$: Base Rate Overuse
$\delta_{s}$ or $\delta_{f}<1$: Base Rate Neglect $\delta_{s}$ or $\delta_{f}<1$: Base Rate Neglect
@@
$\alpha_{0}+\alpha_{Seq}>1$ or $\beta_{0}+\beta_{Seq}>1$: Hot Hand Fallacy
@@
Confidence Biases
$\nu>1$: Overconfidence
$\nu<1$: Underconfidence
@@ 4.3 The effect of each individual bias
In order to facilitate the comprehension and enhance the comparability of biases in Table 1 for the individual analysis, I construct a measure that is informative of the relative importance of each bias. This measure identifies the specific effect of each bias in the relative changes of mean and variance for each subject. In order to achieve this, I analyze the distortions to the Bayesian expected value and Bayesian variance caused by exhibiting a particular kind of bias. The idea is to compare these Bayesian values with a bias-specific measure of how the expected value and variance would look like if the subject exhibited only the specific bias we are interested in. I call these measures Bias-specific expected value and Bias-specific variance. The Bayesian expected value and variance of the posterior beta distribution are given by:
@@
The bias-specific expected value $E_{Bias}$ and bias-specific variance $Var_{Bias}$ [^18] In the case of confidence biases, $Var_{Bias}$ is already given by $\tilde{Var}_{n}$, and $E_{Bias}$ would be unaffected. are given by:
@@
where $a_{bias}$ and $b_{bias}$ are the hypothetical parameters of a posterior beta distribution if they were distorted by only one specific bias. [^19] Because we are looking at the effect of specific biases for each subject, it is theoretically possible that the effect of the bias in isolation is so strong that it makes $a_{bias}$ or $b_{bias}$ negative. This would make $E_{Bias}$ and $Var_{Bias}$ uninterpretable. Therefore, it is assumed that the maximum effect that a bias in isolation can have, is such that the associated parameter $a_{bias}$ or $b_{bias}$ is equal to 0 + $\varepsilon$. It is worth nothing that these cases were very rare, as it would require strongly updating against the information signal. For example, suppose we want to evaluate the impact of overinference. Suppose a particular subject exhibits overinference only in successes: That is $\alpha_{0}>1$. Then, $a_{bias}=a_{Overinfernce}=\alpha_{0}k+a_{0}$ and $b_{bias}=b_{n}$. Equations (19) and (20) for every other bias are calculated analogously.
@@
Equations (21) and (22) allow us to quantify the discrepancy between the Bayesian expected value and variance, and the bias-specific measures. These discrepancies serve as indicative measures of the influence of each bias on the overall inference process. Subsequently, by assessing these differences across subjects, we can assess the relative impact of each individual bias.
@@ 5.1 Population-level analysis
Table 2 compares the baseline model (equations (12) and (13)) with the complete model (equations (14) (15) (16)) at the population level. Columns (1) and (2) of the table refer to the baseline model, while columns (3), (4) and (5) correspond to the complete model. All regressions in table 2 are run with clustered standard errors by participant. [^20] A negligible fraction of observations (4 out of 2640, or 0.1%) yielded slightly negative values for $a_{0}$ and $b_{0}$ due to a minor margin of error in the graphical interface. This only occurred in rare cases where the variance was exceptionally high, and the expected value was near the extremes (0.01 or 0.99). These observations were excluded from the regressions to maintain consistency.
Importantly, the significance of each coefficient does not necessarily show whether an estimate is significantly different from 0, but whether it is significantly different from the corresponding Bayesian value for a given coefficient (see Table 1). In the baseline model, there are two significant biases at the population level: Overinference and base rate neglect. We see overinference both in successes ($\hat{\gamma}_{s}>1$; $\hat{\gamma}_{s}=37.339$ ) and failures ($\hat{\gamma}_{f}>1$; $\hat{\gamma}_{f}=76.119$). This can be interpreted as subjects, on average, interpreting each red ball they observe as 37 red balls; and each blue ball they see as 76 blue balls, when they form their beliefs. And we see base rate neglect (also in successes ($\hat{\delta}_{s}<1$; $\hat{\delta}_{s}=0.017$) and failures $\hat{\delta}_{f}<1$; $\hat{\delta}_{f}=-0.0002$)). This can be interpreted as subjects ignoring their priors (number of red/blue balls they expected to see) when they form their beliefs.
Despite the significance of overinference and base rate neglect at the population level in the baseline model, the effect of overinference seems to vanish when all of the biases of the model are controlled for (Columns (3), (4) and (5)), and base rate neglect is the only remaining significant bias. The complete model also minimizes the AIC and BIC as show in Table 3 (Appendix C). This precisely shows that the lack of inclusion of a wider array of biases within the same model, makes some biases look artificially significant. [^21] Additionally, comparing the adjusted $R^{2}$ column for the individual baseline and complete models in Table 4 further emphasizes this point.
However, neither the baseline nor the complete model seem to be explanatory of subject’s posterior belief formation at the population level. This can be seen by the extremely low goodness of fit ($R^{2}$) of Columns (1) to (4) in Table 2. Yet, this is not the result of subjects being completely noisy, but rather a consequence of individual heterogeneity. That is, subjects are distinctly biased, but they exhibit widely different biases. These distinct biases average out at the population level, but clearly present at the individual level. Table 4 (Appendix C) compares the $R^{2}$ of all the equations at the population with the mean of all $R^{2}$s in these same equations after running the model at the individual level.
Table 2: Baseline and Complete models at the population level. Significance is with respect to Bayesian values. Clustered standard errors.
@@
$a$ $posterior$ $b$ $post$ $a$ $post$ $b$ $post$ $Variance$ $post$
@@
$Failures:preference$-27.602
$Failures:Seq_{neg}$ 102.488
$Confirmation$ 0.001 --0.096
@@
$Bayesian$ $variance$ 0.983
$Constant$ 0.002∗∗
@@
Note: p$<$0.1; ∗∗p$<$0.05; ∗∗∗p$<$0.01 from the Bayesian estimate
As one can see in the first and second column of table 4, the goodness of fit from the population to the individual model shows a substantial improvement. [^22] With the only exception of the variance equation (Eq. (16)). This is also the case for both the AIC and BIC criteria (columns (3) and (4) of table 3), which tremendously diminish when compared to running the models at the population level, with the only exception of the variance equation (equation (16)), which performs better at the population level. This is evidence pointing to the fact that there is considerable individual heterogeneity in the data, and therefore, the model would benefit from an individual analysis.
@@ 5.2 Individual-level analysis
Figures 4(a) and 4(b) provide an overview of the biases detected at the individual level within the context of the two models: the baseline and the complete model. In Figure 4(a), the bar chart depicts the occurrences of statistically significant biases at the 5% level within the baseline model. The color-coded representation distinguishes biases: red means significance solely for successes, blue for failures, and green for significance in both successes and failures. [^23] The column “against” classifies those individuals who update against the information signals. Notably, overinference and base rate overuse emerge prominently. However, Figure 4(b), illustrating individual biases in the complete model, shows a more complete picture.
@@
Here, a broader array of biases is accounted for, with the hot hand fallacy emerging as the most commonly exhibited bias among the subjects. [^24] This is especially surprising given that sequence-related biases, such as the hot hand fallacy, are tested with a reduced number of observations, which generally limits statistical power and thus makes significant findings less expected. Further comparing these models yields intriguing insights. First, at the descriptive level, it becomes evident that individual heterogeneity persists across both models: All tested biases were present among subjects to a certain extent. That is, there is no bias which fully vanishes after incorporating the complete set of biases. Notably, the prevalence of the hot hand fallacy contrasts sharply with the sparsity of confirmatory biases or overconfidence among subjects. Secondly, the complete model makes another important revelation: every subject exhibits at least one form of bias when the full spectrum of biases is considered.
(a) Figure 4a: Number of times a specific bias is found to be significant ($p<0.05$) in the baseline model at the individual level. Overinference and base rate overuse are the most common biases. 15 individuals (17%) exhibit no bias or are too noisy to be distinguished from a Bayesian agent.
@@
(b) Figure 4b: Number of times a specific bias is found to be significant ($p<0.05$) in the complete model at the individual level. There is significant individual heterogeneity. All subjects exhibit at least one type of bias. The hot hand fallacy is found to be the most common bias.
While the “No Bias” column in Figure 4(a) shows that 15 individuals (17% of the sample) could not be categorized, this column is empty in Figure 4(b). This implies that the shift from the baseline to the complete model is able to identify biased individuals, who would otherwise be categorized as either Bayesian or “too noisy”. [^25] Because the individual-level analysis tests many null hypotheses (eleven per subject), one may wonder whether some of these results are driven by a high rate of false positives. To address this concern, section C.2 of the appendix replicates Figures 4(a) and 4(b) correcting for multiple hypothesis testing.
@@
Despite these analyses showing the prevalence of different biases at the individual level, it is also paramount to also consider how much each of these biases drives inference in more tangible and common units. Thus, I apply the methodology outlined in section 4.3, to asses the relative importance of each bias.
@@
Figures 5(a) and 5(b) provide an examination of biases within the complete model, focusing on the expected value deviations from the Bayesian framework for each specific bias. While both figures provide insights into how important these deviations are, they take different approaches. Figure 5(a) presents the expected value deviations for each bias, irrespective of their frequency of occurrence. On the other hand, Figure 5(b) adjusts for frequency by weighting these deviations with the number of times a bias was found to be statistically significant.
@@
(b) Figure 5b: Expected-value deviations adjusted by the amount of times a bias was found to be significant.
In Figure 5(a), two biases, gambler’s fallacy and optimism, stand out with prominent expected value deviations, emphasizing their substantial impact on inference. In contrast, confirmation and disconfirmation bias exhibit less pronounced deviations. Remarkably, even after correcting for frequency in Figure 5(b), gambler’s fallacy and optimism retain their prominence, suggesting their enduring influence on biased inference.
Beyond these observations, Figure 5(a) and 5(b) also reveal interesting patterns. The biases exerting the most substantial influence on expected value deviations cluster into two distinct categories: Motivated beliefs (optimism and pessimism), and biases associated with updating against the information signal, including gambler’s fallacy. Additionally, there is a notable observation regarding the hot hand fallacy: While identified as the most common bias in Figure 4(b), its impact on expected value deviations appears relatively modest in Figure 5(a). However, when it comes to its effect on variance, the hot hand fallacy demonstrates a notably strong influence, as can be seen in Figures 6(a) and 6(b).
Figure 6(a) and 6(b) are analogous to 5(a) and 5(b), but represent variance deviations with respect to the Bayesian framework instead. While the relative contribution of each bias appears to be quite homogeneous (Figure 6(a)), once we correct for the frequency such biases (Figure 6(b)), the hot hand fallacy shows its prominence, while optimism and pessimism come second and third respectively. This underlines the importance of motivated-belief biases in overall inference: Both in expected-value and variance deviations.
@@
(b) Figure 6b: Variance deviations adjusted by the amount of times a bias was found to be significant.
Putting it all together, this subsection emphasizes the importance of incorporating a more complete spectrum of biases into account. When doing so, all biases are present at the individual level and there is no single subject who does not exhibit some kind of bias. It also showcases a key interplay between bias frequency and inference. While, the hot hand fallacy is the most common bias, it is mainly so through its effect on variance. That is, overreaction to a consecutive set of signals in the same direction does not seem to distort the perception of how likely the outcome is as much as distorting how certain individuals are. However, optimism and pessimism seem to be driving inference in both dimensions. Overall, in spite of the presence of individual heterogeneity, two categories of biases (sequence-related biases and motivated-belief biases) clearly stand out above the rest.

### v4

@@ 4.2 Complete model regressions
Variables in equations (14) and (15) are identical to the baseline model with the only exception of variable $c$, the relative measure of confirmation, as described in section 2.2.2. Equation (16) is almost identical to equation (11) in the interpretation of its variables and the $\nu$ parameter indicating over or underprecision ($\nu<1$ vs $\nu>1$). The inclusion of an intercept $\eta$ and an error term $\varepsilon_{v}$ are the only differences. [^18] Similarly to the baseline model, equations (14) and (15) do not have an intercept for the very same reason as outlined in section 4.1.
Parameters $\rho$ and $\delta$ in equations (14) and (15) capture confirmation biases, and base-rate distortions respectively. As in the baseline case, equations (14) and (15) differ by acknowledging that these parameters may differ between successes and failures ($\rho_{s}$ vs $\rho_{f}$) and ($\delta_{s}$ vs $\delta_{f}$). However, the notable difference between equations (14), (15) and (9),(10) lies in accommodating the biases of section 2.2.3. This involves the inclusion of three dummy variables – $I_{Pref}$, $I_{Seq_{s}}$, and $I_{Seq_{f}}$ – interacting with the number of successes (in (14)) or failures (in (15)). $I_{Pref}$ equals 1 when the subject faces a decision with preferences over the state-space (that is, when beliefs about a dollar urn were reported) and 0 otherwise. $I_{Seq_{s}}$ ($I_{Seq_{f}}$) equals 1 when the last three balls observed in the second sequence are red (blue). By doing so, one can distinguish motivated beliefs (optimism and pessimism), sequence-related biases (gambler’s fallacy and hot-hand fallacy) and over/under inference. The presence of those biases is tested by comparing whether the estimated parameters are significantly different from their Bayesian values. Table 1 summarizes the baseline and complete models, and specifically shows the ranges of values of the different parameters, which correspond to each bias.
@@
$\gamma_{s}$ or $\gamma_{f}\in[0,1)$: Underinference $\alpha_{0}$ or $\beta_{0}\in[0,1)$: Underinference
$\gamma_{s}$ or $\gamma_{f}<0$: Against signal $\gamma_{s}$ or $\gamma_{f}<0$: Against signal
Base-Rate Biases
$\delta_{s}$ or $\delta_{f}>1$: Base-Rate Overuse $\delta_{s}$ or $\delta_{f}>1$: Base-Rate Overuse
$\delta_{s}$ or $\delta_{f}<1$: Base-Rate Neglect $\delta_{s}$ or $\delta_{f}<1$: Base-Rate Neglect
@@
$\alpha_{0}+\alpha_{Seq}>1$ or $\beta_{0}+\beta_{Seq}>1$: Hot-Hand Fallacy
@@
Precision Biases
$\nu>1$: Underprecision
$\nu<1$: Overprecision
@@ 4.3 The effect of each individual bias
A bias may be frequently detected across subjects and yet have only a modest quantitative effect on inference when it appears. Conversely, a less prevalent bias may generate large distortions in posterior beliefs. For this reason, beyond classifying which biases are present, I also construct a measure of the relative importance of each bias in driving deviations from the Bayesian benchmark. This measure identifies the specific effect of each bias in the relative changes of mean and variance for each subject. In order to achieve this, I compare the Bayesian expected value and Bayesian variance to bias-specific counterfactual moments that isolate the effect of one particular bias. I refer to these objects as Bias-specific expected value and Bias-specific variance. The Bayesian expected value and variance of the posterior beta distribution are given by:
@@
The bias-specific expected value $E_{Bias}$ and bias-specific variance $Var_{Bias}$ [^19] In the case of precision biases, $Var_{Bias}$ is already given by $\tilde{Var}_{n}$, and $E_{Bias}$ would be unaffected. are given by:
@@
where $a_{bias}$ and $b_{bias}$ are the hypothetical parameters of a posterior beta distribution if they were distorted by only one specific bias. [^20] Because we are looking at the effect of specific biases for each subject, it is theoretically possible that the effect of the bias in isolation is so strong that it makes $a_{bias}$ or $b_{bias}$ negative. This would make $E_{Bias}$ and $Var_{Bias}$ uninterpretable. Therefore, it is assumed that the maximum effect that a bias in isolation can have, is such that the associated parameter $a_{bias}$ or $b_{bias}$ is equal to 0 + $\varepsilon$. It is worth noting that these cases were very rare, as it would require strongly updating against the information signal. For example, suppose we want to evaluate the impact of overinference. Suppose a particular subject exhibits overinference only in successes: That is $\alpha_{0}>1$. Then, $a_{bias}=a_{Overinference}=\alpha_{0}k+a_{0}$ and $b_{bias}=b_{n}$. Equations (19) and (20) for every other bias are calculated analogously.
@@
Equations (21) and (22) allow us to quantify the discrepancy between the Bayesian expected value and variance, and the bias-specific measures. These discrepancies serve as indicative measures of the influence of each bias on the overall inference process.
@@ 5.1 Population-level analysis
Table 2 reports the baseline model (equations (12) and (13)) and the complete model (equations (14), (15) and (16)) estimated at the population level. Columns (1) and (2) correspond to the baseline model, while columns (3), (4) and (5) correspond to the complete model. All regressions in Table 2 are estimated with standard errors clustered at the participant level. [^21] A negligible fraction of observations (4 out of 2640, or 0.1%) yielded slightly negative values for $a_{0}$ and $b_{0}$ due to a minor margin of error in the graphical interface. This only occurred in rare cases where the reported variance was exceptionally high and the expected value was near the extremes (0.01 or 0.99). These observations were excluded from the regressions.
@@
$a$ posteriorposterior $b$ postpost $a$ postpost $b$ postpost VarianceVariance postpost
@@
$Failures:preference$ -27.703
$Failures:Seq_{neg}$ 102.810
ConfirmationConfirmation 0.001 -0.097
@@
BayesianBayesian variancevariance 0.992
ConstantConstant 0.002∗∗
@@
Note: p<<0.1; ∗∗p<<0.05; ∗∗∗p<<0.01 relative to the Bayesian benchmark
Table 2: Baseline and complete models at the population level. Significance is with respect to Bayesian values. Clustered standard errors by participant.
As in the rest of the paper, significance is defined relative to the corresponding Bayesian benchmark rather than relative to zero (see Table 1). In the baseline specification, two biases are significant at the population level: overinference and base-rate neglect. In particular, the coefficients on successes and failures are significantly above their Bayesian benchmarks, while the coefficients on prior beliefs are significantly below one. Thus, when the data are summarized at the aggregate level, the baseline model suggests that subjects overweight the information signal and underweight prior beliefs.
Once the full set of biases is incorporated, the evidence for population-level overinference disappears, whereas base-rate neglect remains significant. This is the main comparison between the baseline and complete specifications at the aggregate level: when a richer set of biases is factored in, baseline overinference estimates appear to be partly absorbing other distortions. This interpretation is reinforced by the information criteria reported in Table 3 in Appendix C, where the complete model improves upon the baseline specification. In addition, Appendix E reports a Grether-style reduced-form specification of the same data in the traditional binary-state framework. The resulting estimates are qualitatively consistent both with the classic belief-updating literature and with the broad aggregate patterns reported here.
At the same time, Table 2 shows that the population-level equations explain little of the variation in posterior mean reports. The goodness of fit is extremely low in columns (1) to (4), for both the baseline and complete specifications. However, this should not be interpreted as the model being unable to explain subjects’ updating behavior. Rather, it reflects the fact that different subjects exhibit different distortions, which tend to average out in the population regression. Consistent with this interpretation, Table 4 in Appendix C shows that the goodness of fit of all equations improves markedly when the model is estimated separately for each participant, while Table 3 shows analogous improvements in AIC and BIC.
These results point to substantial individual heterogeneity in belief-updating distortions and motivate the individual-level analysis that follows.
@@ 5.2 Individual-level analysis
Figures 4(a) and 4(b) provide an overview of the biases detected at the individual level within the context of the two models: the baseline and the complete model. In Figure 4(a), the bar chart depicts the occurrences of statistically significant biases at the 5% level within the baseline model. The color-coded representation distinguishes biases: red means significance solely for successes, blue for failures, and green for significance in both successes and failures. Notably, overinference and base-rate overuse emerge prominently. However, Figure 4(b), illustrating individual biases in the complete model, shows a more complete picture.
@@
Here, a broader array of biases is accounted for, with the hot-hand fallacy being the most commonly exhibited bias among the subjects. [^22] This is especially surprising given that sequence-related biases, such as the hot-hand fallacy, are tested with a reduced number of observations, which generally limits statistical power and thus makes significant findings less expected. Further comparing these models yields interesting insights. First, at the descriptive level, it becomes evident that individual heterogeneity persists across both models: All tested biases were present among subjects to a certain extent. That is, there is no bias which fully vanishes after incorporating the complete set of biases. Notably, the prevalence of the hot-hand fallacy contrasts sharply with the sparsity of confirmatory biases or underprecision among subjects. Second, moving from the baseline to the complete model substantially reduces the set of subjects who remain uncategorized.
@@
(a) Figure 4a: Number of times a specific bias is found to be significant ($p<0.05$) in the baseline model at the individual level.
(b) Figure 4b: Number of times a specific bias is found to be significant ($p<0.05$) in the complete model at the individual level.
While the “No Bias” bar in Figure 4(a) shows that 15 individuals (17% of the sample) could not be categorized –that is, they are either too noisy to classify or statistically indistinguishable from Bayesian updaters– in the complete model (Figure 4(b)) this number falls to 2. Since the complete model tests a larger number of candidate distortions, Appendix D.2 revisits these comparisons under a model-specific multiple-hypothesis testing correction. The main conclusions remain largely unchanged: all biases continue to be present, the hot-hand fallacy is the most commonly exhibited bias, and the complete model still classifies substantially more subjects than the baseline model.
@@
Despite these analyses showing the prevalence of different biases at the individual level, this does not tell us which biases matter most for inference. A bias may be common but quantitatively modest, or relatively less frequent but highly consequential when it appears. Thus, I apply the methodology outlined in section 4.3, to assess the relative importance of each bias.
@@
Figures 5(a) and 5(b) examine biases within the complete model, focusing on the expected-value deviations from the Bayesian framework for each specific bias. While both figures provide insights into how important these deviations are, they take different approaches. Figure 5(a) presents the expected value deviations for each bias, irrespective of their frequency of occurrence. On the other hand, Figure 5(b) adjusts for frequency by weighting these deviations with the number of times a bias was found to be statistically significant.
@@
(b) Figure 5b: Expected-value deviations adjusted by bias prevalence (at $p<0.05)$.
In Figure 5(a), two biases, gambler’s fallacy and optimism, stand out with prominent expected value deviations ($\approx 0.38$ and $\approx 0.35$ distance units to the Bayesian posterior expected value respectively), which emphasizes their substantial impact on inference. In contrast, confirmation and disconfirmation bias exhibit less pronounced deviations. Remarkably, even after correcting for frequency in Figure 5(b), gambler’s fallacy and optimism retain their prominence, suggesting a clear influence on biased inference. [^23] Figures 5(b) and 6(b) should be interpreted ordinally.
Beyond these observations, Figure 5(a) and 5(b) also reveal interesting patterns. The biases exerting the most substantial influence on expected-value deviations cluster into two distinct categories: Motivated beliefs (optimism and pessimism), and biases associated with updating against the information signal, including gambler’s fallacy. Additionally, there is a notable observation regarding the hot-hand fallacy: While identified as the most common bias in Figure 4(b), its impact on expected value deviations appears relatively modest in Figure 5(a). However, when it comes to its effect on variance, the hot-hand fallacy demonstrates a notably strong influence, as can be seen in Figures 6(a) and 6(b).
Figure 6(a) and 6(b) are analogous to 5(a) and 5(b), but represent variance deviations with respect to the Bayesian framework instead. While the relative contribution of each bias appears to be quite homogeneous (Figure 6(a)), once we adjust for bias prevalence such biases (Figure 6(b)), the hot-hand fallacy shows its prominence, while optimism and pessimism come second and third respectively. This underlines the importance of motivated-belief biases in overall inference: Both in expected-value and variance deviations.
@@
(b) Figure 6b: Variance deviations adjusted by bias prevalence (at $p<0.05)$.
Putting it all together, the individual-level analysis shows that the complete model provides a richer account of heterogeneity than the baseline model, and identifies sequence-related and motivated-belief biases as the main sources of departures from Bayesian updating.
5.3 Co-occurrence patterns across belief biases
Biases rarely appear in isolation. To summarize heterogeneity in a compact and easily interpretable way, I study how the biases estimated in the complete model co-occur across subjects. [^24] This analysis is descriptive and exploratory: it was not pre-registered in this exact form, and it is included as a complement to the individual-level estimates. Let $j\in\{1,\dots,J\}$ index the set of bias labels considered in the complete model (e.g., over-/under-inference, base-rate distortions, motivated-belief asymmetries, sequence-related biases, (over/under)precision, and confirmation/disconfirmation). For each subject $i\in\{1,\dots,N\}$, I define a binary indicator $B_{ij}\in\{0,1\},$ which equals one if subject $i$ is classified as exhibiting bias $j$ (according to the individual-level estimation), and zero otherwise. Stacking these indicators yields an $N\times J$ subject-by-bias matrix $B$.
For each bias $j$, I compute its prevalence
$\hat{\upsilon}_{j}\;=\;\frac{1}{N}\sum_{i=1}^{N}B_{ij}$ (23)
,which is simply a subject-normalized score for Figure 4(b) (e.g. hot-hand: $\hat{\upsilon}\approx 0.44$, overinference $\hat{\upsilon}\approx 0.31$, base-rate distortions $\hat{\upsilon}\approx 0.29$ etc.). [^25] Because some biases are low-prevalence, associations involving these biases should be interpreted cautiously.
To quantify co-occurrence between two binary bias indicators $B_{\cdot j}$ and $B_{\cdot j}$, I compute the phi correlation coefficient, $\phi_{jj^{\prime}}=\mathrm{corr}(B_{\cdot j},B_{\cdot k}),$ which coincides with the Pearson correlation computed on $\{0,1\}$-valued variables. Positive values indicate that the two biases tend to be jointly exhibited, while negative values indicate that they tend to be mutually exclusive.
Figure 7: Bias co-occurrence heatmap. Each cell reports the phi correlation $\phi_{jj^{\prime}}=\mathrm{corr}(B_{\cdot j},B_{\cdot j^{\prime}})$ between two bias indicators across subjects. Biases are ordered by hierarchical clustering (average linkage) using distance $1-\phi_{jj^{\prime}}$.
I use two complementary visualizations. Figure 7 displays the full matrix of pairwise co-occurrence coefficients. The hierarchical ordering is used purely for readability, so that positively related biases appear as reddish blocks and systematic negative relationships as bluish contrasts. The dendrogram along the margins shows the resulting clustering.
Figure 8: Bias co-occurrence network. Nodes represent biases and node size is proportional to prevalence $\hat{\upsilon}_{j}$. A red edge is drawn when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=1),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=1)\}\geq 0.5$. A blue edge is drawn when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=0),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=0)\}\geq 0.5$. Edge thickness is proportional to $|\phi_{jj^{\prime}}|$.
Figure 8 provides a condensed summary of the strongest co-occurrence patterns, as a network between biases. Nodes represent biases and node size is proportional to prevalence $\hat{\upsilon}_{j}$. An edge between two biases $j$ and $j^{\prime}$ is drawn when their association is strong enough to have a transparent probabilistic interpretation. Specifically, I draw a red edge (positive co-occurrence) when observing one bias makes the other more likely than not, i.e. when $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=1),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=1)\}\geq 0.5$. Conversely, I draw a blue edge (negative co-occurrence) when the reverse holds in the sense of near-mutual exclusivity, i.e. when the presence of bias $j$ makes it more unlikely that bias $j^{\prime}$ is happening ( $\max\{\Pr(B_{ij}=1\mid B_{ij^{\prime}}=0),\,\Pr(B_{ij^{\prime}}=1\mid B_{ij}=0)\}\geq 0.5$). Edge thickness is proportional to $|\phi_{jj^{\prime}}|$.
Both figures point to a small number of clear qualitative patterns. First, the two sequence-related biases exhibit a strong negative relationship: hot-hand and gambler’s fallacy tend not to be jointly exhibited. In the individual-level data, $\phi(\text{Gambler},\text{HotHand})\approx-0.36$, and the conditional probabilities reflect this asymmetry: $\Pr(\text{HotHand}=1\mid\text{Gambler}=1)\approx 0.11$ while $\Pr(\text{HotHand}=1\mid\text{Gambler}=0)\approx 0.54$. This relationship appears as a blue contrast in the heatmap and as the only prominent blue edge connecting the two sequence nodes in the network. [^26] Note that even though this strong negative correlation is highly intuitive, these biases are (in theory) not necessarily mutually exclusive by definition, as there could be a success-failure asymmetry.
Second, and perhaps more importantly, this divide permeates through the rest of the bias set. The resulting graph from Figure 8 –which highlights the strongest co-occurrence relationships– has a roughly path-like form, with the sequence-related biases (hot-hand and gambler’s fallacy) acting as a hub that separates two main network components. On one side, prior-based distortions and updating against the signal are closely connected, with the gambler/against link and its connections to base-rate overuse forming a compact cluster. On the other side, a set of “jumping-to-conclusions” distortions appear jointly: overprecision and optimism are directly connected with the hot-hand fallacy, while overinference, pessimism and motivated belief asymmetries are only strongly connected to the hot-hand effect through overprecision. The same component also connects to confirmation bias. [^27] The presence of confirmation, and underprecision in this component should be interpreted with caution given their relatively low prevalence, but the position of confirmation bias in the network is consistent with appearing alongside strong updating and concentrated posterior beliefs.
Finally, several pairs of biases that correspond to opposing distortions are negatively related in the co-occurrence matrix. Beyond the hot-hand/gambler opposition discussed above, base-rate neglect and base-rate overuse are negatively related, $\phi(\text{BRN},\text{BRO})\approx-0.24$, and overinference is negatively related to underinference, $\phi(\text{Overinfer},\text{Underinfer})\approx-0.21$ (etc.) At the same time, some conceptually related biases co-occur strongly: e.g. gambler’s fallacy is positively associated with updating against the signal ($\phi\approx 0.36$), and the hot hand with overprecision ($\phi\approx 0.17$).
Overall, the co-occurrence structure suggests that biases cluster into a small number of recurring patterns rather than appearing independently, and further highlights the importance of sequence-related biases. Because this analysis is partly exploratory and some biases are low-prevalence, I treat these patterns as descriptive regularities rather than as definitive typologies.